OpenStudy (anonymous):

if y=-x and y is not = to 0 what is the value of x^2013/y^2013?

OpenStudy (anonymous):

\[\frac{x^{2013}}{y^{2013}}=[\frac{x}y]^{2013}=[\frac{x}{-x}]^{2013}=(-1)^{2013}\\ \text{Since 2012 is odd, we know }(-1)^{2013}\text{ must therefore be }-1.\]

OpenStudy (anonymous):

2013 **

OpenStudy (anonymous):

ok thank you...i have more problems...

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

r u there?